OpenStudy (bbb911):

complex fractions (3/4 Divided by 6/5) - (3/4 X 6/5)

OpenStudy (anonymous):

-11/40

OpenStudy (bbb911):

kk do you mind showing me how its done i kept trying to get that answer and no use

OpenStudy (anonymous):

= {(3/4)*(5/6)}-{(3/4)*(6/5)} = (15/24)-(18/20) = (75-108)/120 = -33/120 = -11/40